Whispering has the following inventory data:

July   1 Beginning inventory 46 units at $110
5 Purchases 276 units at $103
14 Sale 184 units
21 Purchases 138 units at $106
30 Sale 129 units


Assuming that a periodic inventory system is used, what is the cost of goods sold on a FIFO basis?


$15463
$15555
$32653
$32561
Answer #1
Units sold 313 =184+129
Cost of goods sold :
From Beginning inventory 5060 =46*110
From July 5 purchases 27501 =(313-46)*103
Cost of goods sold 32561
Option D $32561 is correct